(5 points) A milk processor monitors the number of bacteria per milliliter in raw milk received at the factory. A random sample of 10 one-milliliter specimens of milk supplied by one producer gives the following data (assume the population is Normal):

5370

4890

5100

4500

5260

5150

4900

4760

4700

4870

Construct and interpret a 90% confidence interval for the population mean μ.
Answer #1

calculate mean and standard deviation for the given sample data

Count, N: 10 Sum, Ex: 49500 Mean, X 4950 Variance, s2: 72066.666666667 Steps N 1 S = (di - T)? N-1 i=1 52 - E(X; - 8)2 N-1 (5

The number of degrees of freedom are df = 10 – 1=9, and the significance level is a = 0.1. Based on the provided information,

What this means is if we repeated experiments, the proportion of such intervals that contain u would be 90%


answered by: ANURANJAN SARSAM
Answer #2

calculate mean and standard deviation for the given sample data

Count, N: 10 Sum, Ex: 49500 Mean, X 4950 Variance, s2: 72066.666666667 Steps N 1 S = (di - T)? N-1 i=1 52 - E(X; - 8)2 N-1 (5

The number of degrees of freedom are df = 10 – 1=9, and the significance level is a = 0.1. Based on the provided information,

What this means is if we repeated experiments, the proportion of such intervals that contain u would be 90%